OpenStudy (anonymous):

Just want to check my answer :) Pyramid A is a square pyramid with a base side length of 14 inches and a height of 6 inches. Pyramid B has a volume of 3136 cubic inches. How many times bigger is the volume of pyramid B than pyramid A? Give your answer as a percent. Provide an explanation and proof for your answer to receive full credit.

OpenStudy (anonymous):

my answer - volume of sq pyramid=1/3 b^2*h volume of pyramid A = 1/3 * 196 * 6 =196*2 =392 Pyramid B has a volume of 3136 cubic inches vol...a/b =3136/392=8times percent change = 392/3136 * 100 =12.5%

OpenStudy (kaederfds):

i think that's true

OpenStudy (anonymous):

this is kinda ambiguous. It's asking *how many times* bigger.

OpenStudy (anonymous):

So i don't think it's asking for percent increase

OpenStudy (anonymous):

So i am wrong? just wondering?

OpenStudy (anonymous):

I would go for 8 times bigger.

OpenStudy (anonymous):

But it is asking for a percent :) so would it be 12.5% or 800%

OpenStudy (anonymous):

well, 8 = 800%

OpenStudy (anonymous):

percentage can mean different things depends on the context. so 800% doesn't mean 800% more but rather 800% of the volume of A

OpenStudy (anonymous):

lol so I am just going to switch it at the end to 3136/392 * 100 =800 ? right?

OpenStudy (anonymous):

you probably meant 3136/392 = 800%

OpenStudy (anonymous):

okay :P Thank you!!! @Kaederfds @xxferrocixx